# Chemical Sciences Division, Lawrence Berkeley National Laboratory, Berkeley, CA 94720, USA

### v1

@@ 3.1 Integration of MGCDB84 and GMTKN55
Retained: X40, XB51 (XB20), XB18 (XB8) (MGCDB84); Excluded: HAL59 (GMTKN55).
@@ 3.2 Reference Updates and Data set adjustments
S22 and S66 are updated with the latest reference values.
@@
The geometry files of MGCDB84 and GMTKN55 were downloaded from the Github repository of ACCDB 168, and we fixed some typos.
@@ 3.3 Newly Added Data Sets
MGCDB84 and GMTKN55 only involve the energetics of main-group element systems, lacking coverage of transition metal systems and various molecular properties. To address these limitations, the GSCDB138 database includes new data sets covering transition metals, main-group metal clusters, electric response properties, vibrational frequencies, and three-body noncovalent interactions. Additionally, new data sets are incorporated to enhance coverage of electron affinities, ionization potentials, and reaction barrier heights. Note that for some data sets, the original papers do not report non-relativistic energies or individual relativistic corrections, making it infeasible to exclude relativistic contributions. Considering that the relativistic effect is usually small for light main-group elements and can be partially and implicitly described for heavy elements by the use of the PP basis sets, these data sets are retained but need to be used cautiously.
@@
3d4dIPSS includes spin states and ionization potentials (IP) for first-row (3d) and second-row (4d) transition metal atoms. For each spin multiplicity, only the lowest-energy state is retained. Theoretical values are adopted from Refs. 154, 155 when available; otherwise, experimental data from Refs. 156, 157 are used after subtracting the Douglas-Kroll-Hess (DKH) correction calculated at the CCSD(T)/aug-cc-pwCVQZ level. Note that many spin multiplicities in Ref. 156 (4dIPSS set) are incorrect, and we have corrected them in our new 3d4dIPSS set.
@@
Pol130 includes static polarizabilities (with directional components) for 130 small molecules. The FD step size for Na is set to 0.003 a.u. to reliably obtain its large value; all other species’ step sizes match the values reported in the original paper. We validated FD results against analytical static polarizabilities in Ref. 49 and removed the data points with large FD errors.
@@
OEEF reports changes in electronic energies under oriented external electric fields relative to the field-free condition. The reference data are calculated at the CCSD(T)/aug-cc-pVQZ level. We observed that B97M-V/aug-pc-3 exhibits large errors for several cases due to artificial spin-symmetry breaking under very strong fields. Notably, this issue disappears when using the def2-QZVPPD basis. The three affected difficult data points are separated into a dedicated OEEFD set for future investigation and are not included in the GSCDB138 benchmark.
@@
V30 contains the vibrational frequencies for molecular dimers across different polarity combinations (polar–polar, polar–nonpolar, and nonpolar–nonpolar). Reference frequencies are calculated using atomic masses of the most abundant isotopes. For numerical stability, the FD step size is adjusted based on frequency magnitude: 0.001 times the normal mode for frequencies over 1000 , 0.003 times the normal mode for frequencies between 1000 and 100 , and 0.01 times the normal mode for frequencies below 100 . We validated FD results from B97M‑V against analytical frequencies169 and excluded any modes where FD errors exceeded 5 and contributed at least 25% of the functional error. As a result, the final set contains no vibrational modes below 40 .
@@
P34 includes AE, IP, and electron affinities (EAs) of systems containing heavy p-block elements up to xenon, further increasing the element coverage of our new database. (The scalar-relativistic correction is not excluded.)
@@
IP30 contains vertical ionization potentials, assembled by selecting the most accurate data points from Ref. 146 and removing overlaps with IP23.
@@ 3.4 Spin Contamination Resolution
To ensure the reliability of the reference values, we address potential issues related to spin contamination. We begin by performing internal stability analyses 170 using the B97X-V functional with a small basis set to determine whether the SCF procedure converges to the true lowest-energy solution in the first run of the unrestricted SCF calculation. Molecules that fail this test are labeled as “Difficult.”
These “Difficult” molecules often exhibit spin contamination. To evaluate whether spin symmetry breaking is physically justified (i.e., essential to describe part of multi-reference effects), we apply the -OOMP2 method with 53, 55. Molecules showing essential spin symmetry breaking are identified by comparing the expected and computed values, as shown in the table of Supporting Information.
@@
Molecule label name | input S | Expected | OOMP2 |
Only use RCCSD(T) level | |||
16_Rxn6_TS_0cat_WCPT27 | 0.0 | 0.0 | 0.37442 |
48_Isomer38_Styrene45 | 0.0 | 0.0 | 0.19079 |
BH9_05_10TS | 0.0 | 0.0 | 0.25343 |
BH9_05_12TS | 0.0 | 0.0 | 0.23487 |
BH9_05_15TS | 0.0 | 0.0 | 0.21359 |
BH9_05_28TS | 0.0 | 0.0 | 0.13999 |
BH9_05_29TS | 0.0 | 0.0 | 0.10417 |
BH9_05_37TS | 0.0 | 0.0 | 0.27310 |
C60ISO_6 | 0.0 | 0.0 | 0.45364 |
C60ISO_8 | 0.0 | 0.0 | 1.00544 |
CARBHB12_1CL | 0.0 | 0.0 | 0.46411 |
CARBHB12_1CL_B | 0.0 | 0.0 | 0.63986 |
CARBHB12_1N | 0.0 | 0.0 | 0.56707 |
CARBHB12_1N_B | 0.0 | 0.0 | 0.60702 |
CARBHB12_1O | 0.0 | 0.0 | 0.51133 |
CARBHB12_1O_B | 0.0 | 0.0 | 0.61909 |
CARBHB12_2CL_B | 0.0 | 0.0 | 0.03752 |
CUAGAU2_M025_Cu6-i0 | 0.0 | 0.0 | 1.04834 |
CUAGAU2_M026_Ag6-i0 | 0.0 | 0.0 | 1.03938 |
CUAGAU2_M027_Au6-i0 | 0.0 | 0.0 | 1.04197 |
CUAGAU2_M028_Cu3-c1-i0 | 0.0 | 0.0 | 0.78269 |
CUAGAU2_M029_Ag3-c1-i0 | 0.0 | 0.0 | 0.71083 |
CUAGAU2_M030_Au3-c1-i0 | 0.0 | 0.0 | 0.76392 |
CUAGAU2_M043_Cu6-c2-i0 | 0.0 | 0.0 | 1.04588 |
CUAGAU2_M044_Ag6-c2-i0 | 0.0 | 0.0 | 1.04020 |
CUAGAU2_M045_Ag6-c2-i0 | 0.0 | 0.0 | 1.04070 |
CUAGAU2_M053_Ag6-i1 | 0.0 | 0.0 | 1.03336 |
CUAGAU2_M054_Au6-i1 | 0.0 | 0.0 | 1.03795 |
CUAGAU2_M058_Cu5-c1-i1 | 0.0 | 0.0 | 1.07203 |
CUAGAU2_M059_Ag5-c1-i1 | 0.0 | 0.0 | 1.05372 |
CUAGAU2_M060_Au5-c1-i1 | 0.0 | 0.0 | 1.05964 |
CUAGAU2_M068_Ag6-c2-i1 | 0.0 | 0.0 | 1.05610 |
CUAGAU2_M079_Cu4-c2-O-i0 | 0.0 | 0.0 | 1.01259 |
CUAGAU2_M080_Ag4-c2-O-i0 | 0.0 | 0.0 | 1.00826 |
CUAGAU2_M118_Cu6-c2-NH3 | 0.0 | 0.0 | 1.04375 |
CUAGAU2_M120_Au6-c2-NH3 | 0.0 | 0.0 | 1.04089 |
CUAGAU2_M124_Cu6-c2-CO | 0.0 | 0.0 | 1.04437 |
CUAGAU2_M125_Ag6-c2-CO | 0.0 | 0.0 | 1.03938 |
CUAGAU_Ag03C1_LIN | 0.0 | 0.0 | 0.71083 |
CUAGAU_Au03C1_LIN | 0.0 | 0.0 | 0.76392 |
CUAGAU_Cu03C1_LIN | 0.0 | 0.0 | 0.78269 |
DIPCS10_h2s_2+ | 0.0 | 0.0 | 0.15004 |
G2RC_6 | 0.0 | 0.0 | 0.58702 |
INV24_PCl3_TS | 0.0 | 0.0 | 1.00683 |
ISOL24_i11p | 0.0 | 0.0 | 0.73205 |
MX35_1208_MgO_01 | 0.0 | 0.0 | 1.00613 |
MME55_CDO_1C | 0.0 | 0.0 | 0.21856 |
MME55_Hc_1_1 | 0.0 | 0.0 | 1.00476 |
MME55_Hc_1_2a | 0.0 | 0.0 | 0.27781 |
UCCSD(T) level | |||
Dip152_CH2-s2+ | 0.0 | 0.0 | 0.59668 |
Dip152_CH2-s2- | 0.0 | 0.0 | 0.59765 |
Dip152_HNS1+ | 0.0 | 0.0 | 0.31640 |
Dip152_HNS1- | 0.0 | 0.0 | 0.31648 |
Pol132_HNS | 0.0 | 0.0 | 0.31644 |
Pol132_HNS0+ | 0.0 | 0.0 | 0.36059 |
Pol132_HNS0- | 0.0 | 0.0 | 0.27974 |
Pol132_HNS1+ | 0.0 | 0.0 | 0.30855 |
Pol132_HNS1- | 0.0 | 0.0 | 0.31659 |
Pol132_HNS2+ | 0.0 | 0.0 | 0.31882 |
Pol132_HNS2- | 0.0 | 0.0 | 0.31882 |
beyond CCSD(T) or experimenta | |||
3d4dIPSS_Y_GS+ | 0.0 | 0.0 | 0.15539 |
ALKBDE10_mgo | 0.0 | 0.0 | 1.00610 |
DAPD_Pd2 | 0.0 | 0.0 | 0.86952 |
IP23_BN | 0.5 | 0.8 | 1.43082 |
IP23_BN_n | 0.0 | 0.0 | 1.02580 |
IP23_C2 | 0.5 | 0.8 | 1.69754 |
IP23_C2_n | 0.0 | 0.0 | 0.89473 |
ORBH36_CO3B_S01 | 0.0 | 0.0 | 0.57909 |
ORBH36_CO3B_S03T | 0.0 | 0.0 | 0.86465 |
W4-17_b2 | 1.0 | 2.0 | 2.81628 |
W4-17_bn | 0.0 | 0.0 | 1.02634 |
W4-17_c2 | 0.0 | 0.0 | 0.87228 |
W4-17_ch2-sing | 0.0 | 0.0 | 0.59060 |
W4-17_cloo | 0.5 | 0.8 | 1.25227 |
W4-17_s3 | 0.0 | 0.0 | 0.12031 |
W4-17_s4-c2v | 0.0 | 0.0 | 0.75903 |
End of Table |
@@ 4 DFT Benchmark Result
To evaluate the performance of modern density functionals, we conduct a comprehensive benchmark using the GSCDB138 database.
@@ 4.1 Functional Candidates and Evaluation Metric
A total of 29 functionals are selected, spanning all rungs of Jacob’s ladder, including the local spin density approximations (LDA), generalized gradient approximations (GGAs), meta-GGAs, hybrid GGAs, hybrid meta-GGAs, and double hybrids (DH). A functional is included if it satisfies at least one of the following criteria: it is widely adopted in the quantum chemistry community, or it ranks among the top two performers in at least one energy category in previous benchmarks such as MGCDB84 or GMTKN55. There are two exceptions. One is the LDA category, for which we include only SPW92, as the MGCDB84 benchmark indicates negligible differences between different LDA parameterizations. The other is DH, where the calculations are very expensive, and we only choose B97M(2) and DSD-PBE-D4 as representative semi-empirical and non-empirical functionals separately.
@@
If a functional does not come with its own dispersion correction, we add D4 201, 202. If no D4 parameter set is available, D3 is used instead. Dispersion energies are computed using the dftd4 203 and simple-dftd3 204 packages. All dispersion parameters are set to their default values as defined in the respective packages (note that using the recent smooth D3S and D4S versions205, 206 of D3 and D4 is expected to make no chemically significant difference to the results205, 206).
All the DFT calculations are performed using a development version of Q-Chem 6.0 207. For most calculations, a (99,590) grid (99 radial shells with 590 grid points per shell) is used for semi-local functional integrals, and SG-1, a subset of (50, 194), is used for non-local VV10 correlation208. A denser grid, up to (500,974) for semi-local integral and (75,302) for non-local integral, will also be used when necessary, such as for dispersion-bonded sets (RG10N, He3), for sets with very large reference values (AE11, AE18), or for some difficult molecules where some functionals strive to get SCF converged in the small grid setup.
We assess the accuracy of each density functional using the mean absolute error (MAE) computed for each data set in the GSCDB138 database, with the exception of Dip146, Pol132, HR46, T144, OEEF, OEEFD, TMD10, MOR13, and TMB11. For Pol132, HR46, T144, OEEF, and OEEFD, mean absolute relative error (MARE) is used instead to account for the wide variation in the magnitudes of the reference values. For Dip146, the mean absolute regularized error defined in the original paper is used 48. For the TMD10, MOR13, and TMB11 sets, we follow the recommendation of the original publications and adopt estimated (weighted) mean absolute deviations.
@@
The performance of each functional is quantified using a normalized error ratio (NER), defined as the ratio between its error and the standard error for a given data set. These ratios are then averaged across all data sets within each property category—barrier heights (BH), electric field responses (EF), vibrational frequencies (FREQ), isomerization energies (ISO), noncovalent interactions (NC), thermochemical properties (TC), and transition metal systems (TM)—as well as across the entire GSCDB138 database to yield an overall mean score for each functional.
@@ 4.2 A Bird’s-Eye View of the Benchmark Results
Figure 1 summarizes the benchmark results for the 29 selected density functionals across seven property categories and their overall mean performance. Overall, functional performance improves along Jacob’s ladder from LDA to double hybrids. The best double hybrid consistently yields the most accurate predictions, followed by the best HMGGA functionals and the best HGGA. In contrast, GGAs and particularly the LDA trail markedly, underscoring the benefits of exact exchange and higher-order correlation. For EF, the best HMGGA (M062X-D3(0)) achieves only a marginally lower mean normalized error (NER) than the best HGGA, B97X-V. Moreover, if root-mean-squared relative/regularized EF errors are used, as in Ref. 48, 49, B97X-V would actually outperform M062X-D3(0). Frequencies are even less sensitive to DFT rung: well-designed MGGAs such as r2SCAN‑D4 and M06L‑D4 perform comparably to HMGGAs, and the best performer overall is PBE0‑D4—–an older, nonempirical HGGA. Nonetheless, hybrid functionals such as B97M-V (except for EF) and B97X-V still achieve strong performance, highlighting their broad transferability.
@@
Among HMGGAs, B97M-V delivers the lowest overall mean NER (1.08) and leads BH, FREQ, ISO, and NC while remaining competitive in other categories except for EF. The second-best performer in this rung is CF22D (overall mean NER is 1.25). It shows broad improvements over earlier Minnesota designs such as MN15-D3(BJ), likely because its training set includes more diverse NC and ISO data. At the same time, B97M-V has weaker EF accuracy compared to B97X-V, and CF22D has poorer performance for EF and FREQ compared to MN15-D3(BJ). This illustrates that HMGGAs can overfit to their training sets and as a consequence, can lose accuracy on properties outside that domain. We therefore recommend B97M-V as the default HMGGA, with CF22D as a somewhat poorer alternative for codes that lack VV10 support. For EF-related properties, M062X-D3(0) or BMK-D3(BJ) may be preferable.
@@
Among HGGAs, B97X-V shows excellent and balanced performance (mean NER = 1.32), ranking as the third-best hybrid functional overall. It performs particularly well for EF, even outperforming B97M-V. Only SOGGA11X-D3(BJ) surpasses B97X-V significantly in BH. Despite its historical popularity, B3LYP performs poorly across most categories and, consistent with earlier conclusions,10, 16 can no longer recommended beyond geometry optimization and vibrational analysis.
@@
In the MGGA category, B97M-V delivers the best overall performance (NER = 1.77), with especially strong results for NC, ranking just behind B97M-V and B97X-V. Its weakest relative performance is for EF and FREQ, where r2SCAN-D4 is the best functional. GGA functionals consistently yield higher errors, with mean NER typically exceeding three. Among them, revPBE-D4 provides the most balanced performance and is recommended for general use, except for EF and FREQ, where N12-D3(0) and OLYP-D4 are superior, respectively. SPW92, the LDA functional, yields the worst overall performance (NER = 12.15), confirming the well-known limitations of the LDA for chemical applications.
@@
Two further points deserve mention. First, FREQ behaves differently from other categories, raising questions about the representativeness of the single set V30. To probe this, we revisited our earlier benchmark 169 and confirmed that PBE0‑D3(BJ) indeed yields the lowest MAE on V30 and across the broader noncovalent frequency benchmark (within the scope of tested functionals). In contrast, B3LYP‑D3(BJ) performs best on covalent frequencies and overall. Thus, V30 alone may not fully capture the diversity of vibrational properties. However, given its reasonable coverage of intramolecular covalent interactions and the computational cost of larger frequency datasets, V30 remains the most practical single-set choice. Second, r2SCAN-D4 outperforms both revPBE-D4 and PBE-D4 across all categories. However, its hybrid counterpart, r2SCAN0-D4, shows similar mean NER to PBE0-D4 but diverges across individual categories. This suggests that improvements achieved from semi-local functional design advances do not necessarily transfer to their hybrid counterparts. Developing hybrid non-empirical functionals may thus require additional physical constraints and design strategies beyond the simple mixing of Hartree–Fock exchange with a non-empirical base.

### v2

@@ 3.1 Integration of MGCDB84 and GMTKN55
Retained: X40, XB51, and XB18 (XB25) (MGCDB84); Excluded: HAL59 (GMTKN55).
@@ 3.2 Reference Updates and Data set adjustments
ISOL23, S22, and S66 are updated with the latest reference values.
@@
XB51 (20 data points) and XB18 (8 data points) from MGCDB84 are merged, and 3 overlapping data points are removed. The resulting set is designated XB25.
PlatonicHD6, PlatonicID6, and PlatonicIG6 are merged into a single set, PlatonicRE18, as their individual performance is of limited interest in the context of such a large database.
@@
The geometry files of MGCDB84 and GMTKN55 were downloaded from the Github repository of ACCDB 169, and we fixed some typos.
@@ 3.3 Newly Added Data Sets
MGCDB84 and GMTKN55 only involve the energetics of main-group element systems, lacking coverage of transition metal systems and various molecular properties. To address these limitations, the GSCDB137 database includes new data sets covering transition metals, main-group metal clusters, electric response properties, vibrational frequencies, and three-body noncovalent interactions. Additionally, new data sets are incorporated to enhance coverage of electron affinities, ionization potentials, and reaction barrier heights. Note that for some data sets, the original papers do not report non-relativistic energies or individual relativistic corrections, making it infeasible to exclude relativistic contributions. Considering that the relativistic effect is usually small for light main-group elements and can be partially and implicitly described for heavy elements by the use of the PP basis sets, these data sets are retained but need to be used cautiously.
@@
3d4dIPSS includes spin states and ionization potentials (IP) for first-row (3d) and second-row (4d) transition metal atoms. For each spin multiplicity, only the lowest-energy state is retained. Theoretical values are adopted from Refs. 156, 157 when available; otherwise, experimental data from Refs. 158, 159 are used after subtracting the Douglas-Kroll-Hess (DKH) correction calculated at the CCSD(T)/aug-cc-pwCVQZ level. Note that many spin multiplicities in Ref. 158 (4dIPSS set) are incorrect, and we have corrected them in our new 3d4dIPSS set.
@@
Pol130 includes static polarizabilities (with directional components) for 130 small molecules. The FD step size for Na is set to 0.003 a.u. to reliably obtain its large value; all other species’ step sizes match the values reported in the original paper. We validated FD results against analytical static polarizabilities in Ref. 50 and removed the data points with large FD errors.
@@
OEEF reports changes in electronic energies under oriented external electric fields relative to the field-free condition. The reference data are calculated at the CCSD(T)/aug-cc-pVQZ level. We observed that B97M-V/aug-pc-3 exhibits large errors for several cases due to artificial spin-symmetry breaking under very strong fields. Notably, this issue disappears when using the def2-QZVPPD basis. The three affected difficult data points are separated into a dedicated OEEFD set for future investigation and are not included in the GSCDB137 benchmark.
@@
V30 contains the vibrational frequencies for molecular dimers across different polarity combinations (polar–polar, polar–nonpolar, and nonpolar–nonpolar). Reference frequencies are calculated using atomic masses of the most abundant isotopes. For numerical stability, the FD step size is adjusted based on frequency magnitude: 0.001 times the normal mode for frequencies over 1000 , 0.003 times the normal mode for frequencies between 1000 and 100 , and 0.01 times the normal mode for frequencies below 100 . We validated FD results from B97M‑V against analytical frequencies170 and excluded any modes where FD errors exceeded 5 or contributed more than 25% of the functional error. As a result, the final set contains no vibrational modes below 40 .
@@
P34 includes AE, IP, and electron affinities (EAs) of systems containing heavy p-block elements up to xenon, further increasing the element coverage of our new database. (The scalar-relativistic correction is not excluded.) We split this set into P34AE, P34IP, and P34EA.
@@
IP30 contains vertical ionization potentials, assembled by selecting the most accurate data points from Ref. 148 and removing overlaps with IP23.
@@ 3.4 Spin Contamination Resolution
To ensure the reliability of the reference values, we address potential issues related to spin contamination. We begin by performing internal stability analyses 171 using the B97X-V functional with a small basis set to determine whether the SCF procedure converges to the true lowest-energy solution in the first run of the unrestricted SCF calculation. Molecules that fail this test are labeled as “Difficult.”
These “Difficult” molecules often exhibit spin contamination. To evaluate whether spin symmetry breaking is physically justified (i.e., essential to describe part of multi-reference effects), we apply the -OOMP2 method with 54, 56. Molecules showing essential spin symmetry breaking are identified by comparing the expected and computed values, as shown in Table S1.
@@ 3.5 Basis Set Considerations
For practical purposes, def2-QZVPPD may still be used for most of these data sets, but substantial errors may occur for certain systems. Notably, large deviations are observed for the AE11 set and for excited He systems in O24x5 (O24_13, O24x4_49, O24x4_50, O24x4_51, O24x4_52). For instance, the error of B97M-V on O24_13 decreases dramatically from –10.7 to +3.0 kcal mol-1 when the basis set is changed from def2-QZVPPD to d-aug-cc-pV5Z.
@@ 4 DFT Benchmark Result
To evaluate the performance of modern density functionals, we conduct a comprehensive benchmark using the GSCDB137 database.
@@ 4.1 Functional Candidates and Evaluation Metric
A total of 29 functionals are selected, spanning all rungs of Jacob’s ladder, including the local spin density approximations (LDA), generalized gradient approximations (GGAs), meta-GGAs, hybrid GGAs, hybrid meta-GGAs, and double hybrids (DH). A functional is included if it satisfies at least one of the following criteria: it is widely adopted in the quantum chemistry community, or it ranks among the top two performers in at least one energy category in previous benchmarks such as MGCDB84 or GMTKN55. There are two exceptions. One is the LDA category, for which we include only SPW92, as the MGCDB84 benchmark indicates negligible differences between different LDA parameterizations. The other is DH, where the calculations are very expensive, and we only choose two available functionals in Q-Chem, B97M(2) and revDSD-PBEP86-D4, as representative semi-empirical and non-empirical functionals.
@@
If a functional does not come with its own dispersion correction, we add D4 202, 203. If no D4 parameter set is available, D3 is used instead. Dispersion energies are computed using the dftd4 204 and simple-dftd3 205 packages. All dispersion parameters are set to their default values as defined in the respective packages (note that using the recent smooth D3S and D4S versions206, 207 of D3 and D4 is expected to make no chemically significant difference to the results206, 207).
All the DFT calculations are performed using a development version of Q-Chem 6.0 208, except for revDSD-PBEP86-D4, which was evaluated with ORCA 6.1 209. For most calculations, a (99,590) grid (99 radial shells with 590 grid points per shell) is used for semi-local functional integrals, and SG-1, a subset of (50, 194), is used for non-local VV10 correlation210. A denser grid, up to (500,974) for semi-local integral and (75,302) for non-local integral, will also be used when necessary, such as for dispersion-bonded sets (RG10N, He3), for sets with very large reference values (AE11, AE18), or for some difficult molecules where some meta-GGAs strive to get SCF converged in the small grid setup.
We assess the accuracy of each density functional using the mean absolute error (MAE) computed for each data set in the GSCDB137 database, with the exception of Dip146, Pol132, HR46, T144, OEEF, OEEFD, O24, O24x4, TMD10, MOR13, and TMB11. For Pol132, HR46, T144, OEEF, and OEEFD, mean absolute relative error (MARE) is used instead to account for the wide variation in the magnitudes of the reference values. For Dip146, the mean absolute regularized error defined in the original paper is used 49. For the TMD10, MOR13, and TMB11 sets, we follow the recommendation of the original publications and adopt estimated (weighted) mean absolute deviations. For O24 and O24x4, weights are assigned according to the general performance of hybrid functionals on each data point within O24. All weighting factors are provided in the supplemental file DatasetEval.xlsx.
@@
The performance of each functional is quantified using a normalized error ratio (NER), defined as the ratio between its error and the standard error for a given data set. These ratios are then averaged across all data sets within each property category—barrier heights (BH), electric field responses (EF), vibrational frequencies (FREQ), isomerization energies (ISO), noncovalent interactions (NC), intramolecular noncovalent interactions (INC), thermochemical properties (TC), and transition metal systems (TM)—as well as across the entire GSCDB137 database to yield an overall mean score for each functional.
@@ 4.2 A Bird’s-Eye View of the Benchmark Results
Figure 2 summarizes the benchmark results for the 29 selected density functionals across seven property categories and their overall mean performance. Overall, functional performance improves along Jacob’s ladder from LDA to double hybrids. The best double hybrid consistently yields the most accurate predictions, followed by the best HMGGA functionals and the best HGGA. In contrast, GGAs and particularly the LDA trail markedly, underscoring the benefits of exact exchange and higher-order correlation. These overall trends and relative rankings are consistent with those observed for the GMTKN55 database. However, when the new EF, FREQ, and TM categories are examined individually, additional distinctions in functional performance emerge that are not captured by GMTKN55. For EF, the best HMGGA (M062X-D3(0)) achieves only a marginally lower mean normalized error (NER) than the best HGGA, B97X-V. Moreover, if root-mean-squared relative/regularized EF errors are used, as in Ref. 49, 50, B97X-V would actually outperform M062X-D3(0). For vibrational frequencies, Jacob’s ladder generally holds, but the top performers—revDSD-PBEP86-D4 and PBE0-D4—are both nonempirical functionals. Nonetheless, hybrid functionals such as B97M-V (except for EF) and B97X-V still achieve strong performance, highlighting their broad transferability.
@@
Among HMGGAs, B97M-V delivers the lowest overall mean NER (1.08) and leads BH, FREQ, INC, and NC while remaining competitive in other categories except for EF. As shown in Figure S1, it achieves “Good” performance (error smaller than that of the fourth-best hybrid) on 88 of the 137 sets. The second-best performer in this rung is CF22D (overall mean NER = 1.25; “Good” count = 74). CF22D is a supervised machine-learned hybrid meta-GGA with more than fifty adjustable parameters, developed using an active-learning strategy based on the MN15 functional form. It shows broad improvements over earlier Minnesota functionals such as its parent, MN15-D3(BJ), likely because its training set includes more diverse NC and ISO data. At the same time, B97M-V has weaker EF accuracy compared to B97X-V, and CF22D has poorer performance for EF and FREQ compared to MN15-D3(BJ). This illustrates that HMGGAs can overfit to their training sets and, as a consequence, can lose accuracy on properties outside that domain. We therefore recommend B97M-V as the default HMGGA, with CF22D as a somewhat poorer alternative for codes that lack VV10 support. For EF-related properties, M062X-D3(0) or BMK-D3(BJ) may be preferable.
@@
Among HGGAs, B97X-V shows excellent and balanced performance (mean NER = 1.31; “Good” count = 69), ranking as the third-best hybrid functional overall. It performs particularly well for EF, even outperforming B97M-V. Only SOGGA11X-D3(BJ) surpasses B97X-V significantly in BH. Despite its historical popularity, B3LYP performs poorly across most categories and, consistent with earlier conclusions,10, 16 can no longer be recommended beyond geometry optimization and vibrational analysis. Our previous benchmark also indicated that a triple-zeta basis set is required for reliable frequency calculations with B3LYP.170 Therefore, the commonly used B3LYP/6-31G* combination should be phased out.
@@
In the MGGA category, B97M-V delivers the best overall performance (mean NER = 1.75), with especially strong results for NC, ranking just behind B97M-V and B97X-V. Its performance is somewhat weaker for EF and FREQ, where r2SCAN-D4 is the best functional. The recently developed deep learning functional Skala is not tested here because its coefficients have not yet been released when our paper is submitted. However, it is expected to outperform traditional MGGA functionals, as it contains additional non-local information. GGA functionals consistently yield higher errors, with mean NER typically exceeding three. Among them, revPBE-D4 provides the most balanced performance and is recommended for general use, except for EF and FREQ, where N12-D3(0) and OLYP-D4 are superior, respectively. SPW92, the LDA functional, yields the worst overall performance (mean NER = 12.25), confirming the well-known limitations of the LDA for chemical applications.
@@
Two further points deserve mention. First, FREQ behaves differently from other categories, raising questions about the representativeness of the single set V30. To probe this, we revisited our earlier benchmark 170 and confirmed that PBE0‑D3(BJ) indeed yields the lowest MAE on V30 and across the broader noncovalent frequency benchmark (within the scope of tested functionals). In contrast, B3LYP‑D3(BJ) performs best on covalent frequencies and overall. Thus, V30 alone may not fully capture the diversity of vibrational properties. However, given its reasonable coverage of intramolecular covalent interactions and the computational cost of larger frequency datasets, V30 remains the most practical single-set choice. Second, r2SCAN-D4 outperforms both revPBE-D4 and PBE-D4 across all categories. However, its hybrid counterpart, r2SCAN0-D4, shows similar mean NER to PBE0-D4 but diverges across individual categories. This suggests that improvements achieved from semi-local functional design advances do not necessarily transfer to their hybrid counterparts.